# Decentralized Prediction Markets and Sports Books

### v3

@@ 4 Decentralized liquidity pooling
Briefly, let $\ell\in L^{\infty}_{+}\backslash\{\mathbf{0}\}$ be a position that the liquidity provider is willing to sell in exchange for a fraction $\alpha>0$ of the liquidity remaining $\Pi$. In particular, the investor wishes to guarantee that this fraction of liquidity remaining is accurate after she places her own bet, i.e., $x=\alpha(\Pi-x+\ell)$ for some $\alpha>0$. Solving for $x=\frac{\alpha}{1+\alpha}(\Pi+\ell)$, we solve the inverse problem to determine the fraction of the liquidity that is being purchased, i.e., $0=C(\frac{\alpha}{1+\alpha}(\Pi+\ell)-\ell;\Pi)$. In other words, prior to the market opening, the payoff for investors who provide funding to the AMM was simply the proportion of reserves that they provided. In contrast, after the market opens and bets have been placed, this is no longer as straightforward. Liquidity providers are still entitled to some portion of the terminal wealth that remains at the AMM (i.e., a proportion of $\Pi$), but that portion may no longer be a simple proportion $\ell/\Pi$ as this may no longer even be deterministic. Instead, the aforementioned calculation determines the deterministic proportion $\alpha$ that will be assigned to the liquidity provider.
@@
$0=\bar{C}(\alpha^{*}(\Pi,\ell)):=C(\frac{\alpha^{*}(\Pi,\ell)}{1+\alpha^{*}(\Pi,\ell)}\Pi-\frac{1}{1+\alpha^{*}(\Pi,\ell)}\ell;\Pi).$ (4.1)
@@
Notably, the bet $\frac{\alpha^{*}(\Pi,\ell)}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell)$ is made in such a way that it will always rebalance to maintain a payout of $\alpha^{*}(\Pi,\ell)$ fraction of the remaining liquidity as new bets are made. Let $\Pi^{*}:=\frac{1}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell)$ be the effective liquidity after pooling the liquidity. When the next trade occurs, the liquidity provider (through, e.g., a smart contract) needs to simultaneously update their position. That is, when a new bet $x\in L^{\infty}$ is made, the trader will be charged $C(x+y-\alpha\Pi^{*};\Pi^{*})$ where $y\in L^{\infty}_{*}$ is the new holdings for the liquidity provider, i.e., $y=\alpha^{*}(\Pi,\ell)(\Pi^{*}-x-(y-\alpha^{*}(\Pi,\ell)\Pi^{*})+C(x+y-\alpha^{*}(\Pi,\ell)\Pi^{*};\Pi^{*})\mathbf{1})$ so that the liquidity provider maintains the $\alpha^{*}(\Pi,\ell)$ fraction of remaining liquidity after the bet occurs.
@@
$\tilde{C}=C(x+[\alpha^{*}\Pi^{*}-\frac{\alpha^{*}}{1+\alpha^{*}}x+\frac{\alpha^{*}}{1+\alpha^{*}}\tilde{C}\mathbf{1}]-\alpha^{*}\Pi^{*};\Pi^{*})=C(\frac{1}{1+\alpha^{*}}x;\Pi^{*})+\frac{\alpha^{*}}{1+\alpha^{*}}\tilde{C}.$
With this foundation for the interaction of the liquidity provision and trades, we now want to specify $\ell\in L^{\infty}_{+}\backslash\{\mathbf{0}\}$ so that the pricing oracles are kept constant, i.e., $P^{b}(\cdot;\Pi)=P^{b}(\cdot;\frac{1}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell))$ and $P^{a}(\cdot;\Pi)=P^{a}(\cdot;\frac{1}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell))$. Providing liquidity in this way means that pooling does not influence the pricing measure quoted by the market.
@@
Recall $\alpha^{*}(\Pi,\ell)$ is the unique root of $\bar{C}(\alpha)$ as defined in (4.1). It trivially follows from that construction that $\alpha^{*}(\Pi,t\Pi)=t$ is a root of $\bar{C}$. Furthermore, and as a direct consequence, the pricing oracles after the liquidity provision or withdrawal are provided by $P^{b}(\cdot;\frac{1}{1+\alpha^{*}(\Pi,t\Pi)}(\Pi+t\Pi))=P^{b}(\cdot;\Pi)$ and similarly for $P^{a}$. ∎
@@
It remains to prove that providing liquidity reduces the costs for counterparties to buy or sell bets. This is proven in the following lemma by noting that $\Pi^{*}=\Pi$ under the proportional liquidity provision $\ell=t\Pi$.
@@ 5 Fee structure
Conceptually, due to the translativity of the cost function $C$, the AMM is indifferent to any almost sure bet. As such, we propose that fees are only collected on the (positive) random portion $x-\operatorname*{ess\,inf}x$ of the bet $x\in L^{\infty}$. That is, for the fixed fee level $\gamma\geq 0$, the cost of purchasing $x\in L^{\infty}$ in the pool $\Pi\in L^{\infty}_{*}$ is given by
$C_{\gamma}(x;\Pi):=(1+\gamma)C(x-\operatorname*{ess\,inf}x;\Pi)+\operatorname*{ess\,inf}x=(1+\gamma)C(x;\Pi)-\gamma\operatorname*{ess\,inf}x.$ (5.1)
In this way, the AMM collects $\gamma C(x-\operatorname*{ess\,inf}x;\Pi)$ in cash to be paid out to the liquidity providers. This surplus is collected by the AMM poolers in compensation for the liquidity they provide. The no-fee setting corresponds exactly to $\gamma=0$; as will be discussed below, for a well-functioning market, we will want to cap the fees at $\gamma\leq 1$ (see Remark 11). We recall from the original setup in (2.1) that the cost function is only applied to the random part of a bet with the essential infimum being accounted for in the liquidity; as such this fee structure matches the logic applied to AMMs generally. Note that the fees $\gamma$ introduced here are not endogenous to the model, but rather are arbitrarily set. Within the numerical case study of Section 6.1.2, we consider how varying $\gamma$ may alter market dynamics and, therefore also, the total collected fees.
@@
By defining the fees on the (positive) random portion $x-\operatorname*{ess\,inf}x$ of the bet $x\in L^{\infty}$, we eliminate a possible violation of the law of one price. Specifically, the cost of buying $x\in L^{\infty}_{+}$ is identical to the cost of buying the fixed payout of $\operatorname*{ess\,sup}x$ and selling $\operatorname*{ess\,sup}x-x$. Neglecting counterparty risks, the payoff of these bets are functionally identical and, in assessing fees only on the random portion of a bet, so are their costs.
@@
Assume $x\in L^{\infty}$. By construction, $\frac{\partial}{\partial\gamma}C_{\gamma}(x;\Pi)=C(x-\operatorname*{ess\,inf};\Pi)$. By Theorem 3.4(1), this is strictly positive so long as $x\not\in\{c\mathbf{1}\;|\;c\in\mathbb{R}\}$. Furthermore, $C_{\gamma}(c\mathbf{1};\Pi)=c$ for any $c\in\mathbb{R}$ by construction for any $\gamma\geq 0.$
Herein, we assume that the collected fees $\gamma C(x-\operatorname*{ess\,inf}x;\Pi)\geq 0$ are immediately distributed to the liquidity providers. As such, when a bet is placed, the pool still updates from $\Pi\in L^{\infty}_{*}$ to $\Pi-x+C(x;\Pi)\mathbf{1}\in L^{\infty}_{*}$ as in the no-fee setting. By collecting and disbursing fees in this manner, the impact on our discussion of liquidity pooling in Section 4 remains unaffected.
@@
    Increasing utility: $u(\Pi-x+C_{\gamma}(x)\mathbf{1})>u(\Pi)$ for $\gamma>0$ and $x\in L^{\infty}\backslash\{c\mathbf{1}\;|\;c\in\mathbb{R}\}$ with strict inequality becoming an equality if if $\gamma=0$ or $x=c\mathbf{1}$ for some $c\in\mathbb{R}$;
@@
    The bounds follow directly from Theorem 3.4(1) applied to the bet $x-\operatorname*{ess\,inf}x$.
@@ 6.1.1 Deterministic Backtesting
Using this bid-ask spread, we were able to determine the (implied) mid-price by normalizing the ask prices to sum to 1. This price oracle, $P(\mathbb{I}_{\mathrm{KC}})$ as displayed in Figure 1(a), drives our backtesting system. Within this construction we consider 2 utility functions: (i) the expected logarithm utility function as discussed in Example 3.2 and (ii) the Liquid StableSwap utility function ($u(x)=\mathbb{E}[\log(x)]+\lambda\log(\mathbb{E}[x])$ with $\lambda=2$) introduced in [7]. With these LBAMMs, we deduce the trades necessary to exactly replicate this pricing oracle (without fees). Assuming $L>0$ cash reserves was initially available to the AMM, the liquidity available at any time can be determined via
@@ 6.1.2 Stochastic Backtesting
We now aim to introduce a stochastic backtesting framework to the Bookmaker money line data we have collected. In contrast to the deterministic backtest of Section 6.1.1 in which the AMM’s price oracle exactly follows the mid-price of the external sports book, here we assume that the underlying market price follows a Brownian motion, starting from the initial mid-price of the external market. We constrain the price process to remain within Bookmaker’s bid-ask spread through a reflection principle. However, rather than assuming the AMM perfectly tracks this price process, we assume trades occur only to capture arbitrage opportunities between the true (stochastic) price process and the price oracle when accounting for the fees $\gamma$. Due to the profits exhibited in the prior backtest, we focus exclusively on a dynamic version of the Liquid StableSwap AMM herein with $u_{t}(x)=\mathbb{E}^{\mathbb{P}_{t}}[\log(x)]+\lambda\log(\mathbb{E}^{\mathbb{P}_{t}}[x])$ with $\lambda=2$ and $\mathbb{P}_{t}(KC)$ is determined by the (implied) mid-price of Bookmaker money line data at time $t$. The choice of dynamic probability measure is such that the stable region of the AMM tracks the mid-price of the external market and is intended to maximize the fees collected. As far as the authors are aware, such a dynamic AMM has not previously been proposed in either prediction or cryptocurrency markets.
@@
Consider now the optimal fee levels for the liquidity provider. Considering Figure 2, though subtle, it can be seen that the optimal fee level $\gamma$ is decreasing in volatility. That is, as volatility increases, the greater number of transactions dominates the aforementioned tradeoff. Intuitively, higher volatility increases the probability that the AMM’s price sufficiently deviates from the Brownian price path to create arbitrage opportunities. In opposition, and as already noted, higher fees reduce the trading activity by widening the no-arbitrage range. Therefore, when volatility is higher, increased potential betting activity around the Brownian price path dominates the reduction in fees collected from each individual transaction. Furthermore, these optimal fee levels (roughly around $\gamma\approx 1\%$) – which result in significant expected profits of between 1% and 6% return in a 2 week period – come at a much lower bid-ask spread than that quoted by Bookmaker. Consequently, the introduction of the AMM creates a win-win situation: liquidity providers can optimize their expected profits and liquidity takers find a more efficient market in which to transact. By the same arguments, the optimal fees are monotonic with respect to Bookmaker’s bid-ask spread but are always dominated by those external fees. In this way, the optimal AMM fee construction must alway increases market efficiency.

### v4

@@ 4 Decentralized liquidity pooling
Briefly, let $\ell\in L^{\infty}_{+}\backslash\{\mathbf{0}\}$ be a position that the liquidity provider is willing to sell in exchange for a fraction $\alpha>0$ of the liquidity remaining $\Pi$. In particular, the investor wishes to guarantee that this fraction of liquidity remaining is accurate after she places her own bet, i.e., $x=\alpha(\Pi-x+\ell)$ for some $\alpha>0$; recall that $\Pi-x+\ell$ is the liquidity that remains in the AMM after the liquidity provision $\ell$ is added, but after the bet $x$ is placed, where the prior liquidity position is $\Pi$. Solving for $x=\frac{\alpha}{1+\alpha}(\Pi+\ell)$, we solve the inverse problem to determine the fraction of the liquidity that is being purchased, i.e., $0=C\left(\frac{\alpha}{1+\alpha}(\Pi+\ell)-\ell;\Pi\right)$. In other words, prior to the market opening, the payoff for investors who provide funding to the AMM was simply the proportion of reserves that they provided. In contrast, after the market opens and bets have been placed, this is no longer as straightforward. Liquidity providers are still entitled to some portion of the terminal wealth that remains at the AMM (i.e., a proportion of $\Pi$), but that portion may no longer be a simple proportion $\ell/\Pi$ as this may no longer even be deterministic. Instead, the aforementioned calculation determines the deterministic proportion $\alpha$ that will be assigned to the liquidity provider. As will be discussed in greater detail below, the payout from this liquidity position will dynamically adjust along with the liquidity of the AMM as new bets are made; the payoff of this bet is the $\alpha$ fraction of the final liquidity $\bar{\Pi}\in L^{\infty}_{*}$ after all bets are placed. It is this dynamic adjustment of the payout which makes this bet a liquidity provision, i.e., the dynamic adjustment takes the opposite position to any new incoming bet thereby decreasing price impacts and increasing market liquidity.
@@
$0=\bar{C}(\alpha^{*}(\Pi,\ell)):=C\left(\frac{\alpha^{*}(\Pi,\ell)}{1+\alpha^{*}(\Pi,\ell)}\Pi-\frac{1}{1+\alpha^{*}(\Pi,\ell)}\ell;\Pi\right).$ (4.1)
@@
Notably, and as briefly mentioned above, the bet $\frac{\alpha^{*}(\Pi,\ell)}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell)$ is made in such a way that it will always rebalance to maintain a payout of $\alpha^{*}(\Pi,\ell)$ fraction of the remaining liquidity as new bets are made. Let $\Pi^{*}:=\frac{1}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell)$ be the effective liquidity after pooling the liquidity. When the next trade occurs, the liquidity provider (through, e.g., a smart contract) needs to simultaneously update their position. That is, when a new bet $x\in L^{\infty}$ is made, the trader will be charged $C(x+y-\alpha\Pi^{*};\Pi^{*})$ where $y\in L^{\infty}_{*}$ is the new holdings for the liquidity provider, i.e., $y=\alpha^{*}(\Pi,\ell)(\Pi^{*}-x-(y-\alpha^{*}(\Pi,\ell)\Pi^{*})+C(x+y-\alpha^{*}(\Pi,\ell)\Pi^{*};\Pi^{*})\mathbf{1})$ so that the liquidity provider maintains the $\alpha^{*}(\Pi,\ell)$ fraction of remaining liquidity after the bet occurs.
@@
$\tilde{C}=C\left(x+\left(\alpha^{*}\Pi^{*}-\frac{\alpha^{*}}{1+\alpha^{*}}x+\frac{\alpha^{*}}{1+\alpha^{*}}\tilde{C}\mathbf{1}\right)-\alpha^{*}\Pi^{*};\Pi^{*}\right)=C\left(\frac{1}{1+\alpha^{*}}x;\Pi^{*}\right)+\frac{\alpha^{*}}{1+\alpha^{*}}\tilde{C}.$
With this foundation for the interaction of the liquidity provision and trades, we now want to specify $\ell\in L^{\infty}_{+}\backslash\{\mathbf{0}\}$ so that the pricing oracles are kept constant, i.e., $P^{b}(\cdot;\Pi)=P^{b}\Big{(}\cdot;\frac{1}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell)\Big{)}$ and $P^{a}(\cdot;\Pi)=P^{a}\left(\cdot;\frac{1}{1+\alpha^{*}(\Pi,\ell)}(\Pi+\ell)\right)$. Providing liquidity in this way means that pooling does not influence the pricing measure quoted by the market.
@@
Recall $\alpha^{*}(\Pi,\ell)$ is the unique root of $\bar{C}(\alpha)$ as defined in (4.1). It trivially follows from that construction that $\alpha^{*}(\Pi,t\Pi)=t$ is a root of $\bar{C}$. Furthermore, and as a direct consequence, the pricing oracles after the liquidity provision or withdrawal are provided by $P^{b}\left(\cdot;\frac{1}{1+\alpha^{*}(\Pi,t\Pi)}(\Pi+t\Pi)\right)=P^{b}(\cdot;\Pi)$ and similarly for $P^{a}$. ∎
@@
It remains to prove that providing liquidity reduces the costs for counterparties to buy or sell bets. Fundamentally, this is what liquidity provision is intended to accomplish as it decreases the price impacts experienced by any bettor. This is proven in the following lemma by noting that $\Pi^{*}=\Pi$ under the proportional liquidity provision $\ell=t\Pi$.
@@ 5 Fee structure
Conceptually, due to the translativity of the cost function $C$, the AMM is indifferent to any almost sure bet. As such, we propose that fees are only collected on the (positive) random portion $x-\operatorname*{ess\,inf}x\mathbf{1}$ of the bet $x\in L^{\infty}$. That is, for the fixed fee level $\gamma\geq 0$, the cost of purchasing $x\in L^{\infty}$ in the pool $\Pi\in L^{\infty}_{*}$ is given by
$C_{\gamma}(x;\Pi):=(1+\gamma)C(x-\operatorname*{ess\,inf}x\mathbf{1};\Pi)+\operatorname*{ess\,inf}x=(1+\gamma)C(x;\Pi)-\gamma\operatorname*{ess\,inf}x.$ (5.1)
In this way, the AMM collects $\gamma C(x-\operatorname*{ess\,inf}x\mathbf{1};\Pi)$ in cash to be paid out to the liquidity providers. This surplus is collected by the AMM poolers in compensation for the liquidity they provide. The no-fee setting corresponds exactly to $\gamma=0$; as will be discussed below, for a well-functioning market, we will want to cap the fees at $\gamma\leq 1$ (see Remark 10). We recall from the original setup in (2.1) that the cost function is only applied to the random part of a bet with the essential infimum being accounted for in the liquidity; as such this fee structure matches the logic applied to AMMs generally. Note that the fees $\gamma$ introduced here are not endogenous to the model, but rather are arbitrarily set. Within the numerical case study of Section 6.1.2, we consider how varying $\gamma$ may alter market dynamics and, therefore also, the total collected fees.
@@
By defining the fees on the (positive) random portion $x-\operatorname*{ess\,inf}x\mathbf{1}$ of the bet $x\in L^{\infty}$, we eliminate a possible violation of the law of one price. Specifically, the cost of buying $x\in L^{\infty}_{+}$ is identical to the cost of buying the fixed payout of $\operatorname*{ess\,sup}x\mathbf{1}$ and selling $\operatorname*{ess\,sup}x\mathbf{1}-x$. Neglecting counterparty risks, the payoff of these bets are functionally identical and, in assessing fees only on the random portion of a bet, so are their costs.
@@
Assume $x\in L^{\infty}$. By construction, $\frac{\partial}{\partial\gamma}C_{\gamma}(x;\Pi)=C(x-\operatorname*{ess\,inf}x\mathbf{1};\Pi)$. By Theorem 3.4(1), this is strictly positive so long as $x\not\in\{c\mathbf{1}\;|\;c\in\mathbb{R}\}$. Furthermore, $C_{\gamma}(c\mathbf{1};\Pi)=c$ for any $c\in\mathbb{R}$ by construction for any $\gamma\geq 0.$
Herein, we assume that the collected fees $\gamma C(x-\operatorname*{ess\,inf}x\mathbf{1};\Pi)\geq 0$ are immediately distributed to the liquidity providers. As such, when a bet is placed, the pool still updates from $\Pi\in L^{\infty}_{*}$ to $\Pi-x+C(x;\Pi)\mathbf{1}\in L^{\infty}_{*}$ as in the no-fee setting. By collecting and disbursing fees in this manner, the impact on our discussion of liquidity pooling in Section 4 remains unaffected.
@@
    Increasing utility: $u(\Pi-x+C_{\gamma}(x)\mathbf{1})>u(\Pi)$ for $\gamma>0$ and $x\in L^{\infty}\backslash\{c\mathbf{1}\;|\;c\in\mathbb{R}\}$ with strict inequality becoming an equality if $\gamma=0$ or $x=c\mathbf{1}$ for some $c\in\mathbb{R}$;
@@
    The bounds follow directly from Theorem 3.4(1) applied to the bet $x-\operatorname*{ess\,inf}x\mathbf{1}$.
@@ 6.1.1 Deterministic Backtesting
Using this bid-ask spread, we were able to determine the (implied) mid-price by normalizing the ask prices to sum to 1. This price oracle, $P(\mathbb{I}_{\mathrm{KC}})$ as displayed in Figure 1(a), drives our backtesting system. Within this construction we consider 2 utility functions: (i) the expected logarithm utility function as discussed in Example 3.2 and (ii) the Liquid StableSwap utility function ($u(x)=\mathbb{E}[\log(x)]+\lambda\log(\mathbb{E}[x])$ with $\lambda=2$) introduced in [7]. With these LBAMMs, we deduce the trades necessary to exactly replicate the mid-price at Bookmaker (without fees). In doing so, we make no assumptions on trader’s behavior but rather assume that the quoted mid-price would, fundamentally, be reflected in any constructed market. Assuming $L>0$ cash reserves was initially available to the AMM, the liquidity available at any time can be determined via
@@ 6.1.2 Stochastic Backtesting
We now aim to introduce a stochastic backtesting framework to the Bookmaker money line data we have collected. In contrast to the deterministic backtest of Section 6.1.1 in which the AMM’s price oracle exactly follows the mid-price of the external sports book, here we assume that the underlying market price follows a Brownian motion, starting from the initial mid-price of the external market. We constrain the price process to remain within Bookmaker’s bid-ask spread through a reflection principle. However, rather than assuming the AMM perfectly tracks this price process, we assume trades occur only to capture arbitrage opportunities between the true (stochastic) price process and the price oracle when accounting for the fees $\gamma$. In this way, as in Section 6.1.1, we make no assumptions on bettor behavior but rather utilize the available betting line data to construct simple price processes in order to find (potential) arbitrage opportunities.
Due to the profits exhibited in the prior backtest, we focus exclusively on a dynamic version of the Liquid StableSwap AMM herein with $u_{t}(x)=\mathbb{E}^{\mathbb{P}_{t}}[\log(x)]+\lambda\log(\mathbb{E}^{\mathbb{P}_{t}}[x])$ with $\lambda=2$ and $\mathbb{P}_{t}(KC)$ is determined by the (implied) mid-price of Bookmaker money line data at time $t$. The choice of dynamic probability measure is such that the stable region of the AMM tracks the mid-price of the external market and is intended to maximize the fees collected. As far as the authors are aware, such a dynamic AMM has not previously been proposed in either prediction or cryptocurrency markets.
@@
Consider now the optimal fee levels for the liquidity provider. Considering Figure 2, though subtle, it can be seen that the optimal fee level $\gamma$ is decreasing in volatility. That is, as volatility increases, the greater number of transactions dominates the aforementioned tradeoff. Intuitively, higher volatility increases the probability that the AMM’s price sufficiently deviates from the Brownian price path to create arbitrage opportunities. In opposition, and as already noted, higher fees reduce the trading activity by widening the no-arbitrage range. Therefore, when volatility is higher, increased potential betting activity around the Brownian price path dominates the reduction in fees collected from each individual transaction. Furthermore, these optimal fee levels (roughly around $\gamma\approx 1\%$) – which result in significant expected profits of between 1% and 6% return in a 2 week period – come at a much lower bid-ask spread than that quoted by Bookmaker. Consequently, the introduction of the AMM creates a win-win situation: liquidity providers can optimize their expected profits and liquidity takers find a more efficient market in which to transact. By the same arguments, the optimal fees are monotonic with respect to Bookmaker’s bid-ask spread but are always dominated by those external fees. In this way, the optimal AMM fee construction must always increases market efficiency.
@@ 6.2 Financial Derivatives
Herein we consider a derivatives payoff as a bet on a market outcome. The LBAMM then represents a type of utility indifference pricing, see, e.g., [10]. In our case, an LBAMM is indifferent between two scenarios – one in which no bet is placed, and the other in which a bet is placed with the liquidity increased by the bet’s price. However, in contrast to a typical derivatives market, we do not consider the market for the underlying asset(s), i.e., it is not possible to hedge using the underlying securities. Instead, the initial liquidity is used to compensate for the fact that there is no hedge possible for the market makers; this initial liquidity is used to guarantee that the LBAMM can pay off the sold derivatives at maturity. That is, within this framework without the underlying market for hedging, options and other derivatives can be viewed as classical bets on market outcomes.
@@
While put options naturally have uniformly bounded payoffs, that is not true for all options. For instance, call options have unbounded payoff and, thus, do not fit into the $L^{\infty}$ framework proposed within this work. However, much as in the risk measurement literature (see, e.g., [17]), call options and other derivatives with unbounded payouts can be adapted to this framework by imposing a cap to the bet, e.g., the adapted call option would take value $x_{T}:=\min\{(S-K)^{+}\,,\,T\}$ for underlying price $S$, strike price $K$, and some cap $T>0$ sufficiently large. Figure 4(a) displays the cost of 100 at-the-money call options with stirke \$1 assuming an initial lognormal distribution; notably, if $T$ is set too high, e.g., so that $T>\operatorname*{ess\,sup}\Pi$, then the cost of these options will be proportional to $T$ to guarantee the positivity of the resulting position $\Pi-x_{T}+C(x_{T};\Pi)\mathbf{1}\in L^{\infty}_{*}$. However, if we scale the initial liquidity with the cap $T$ then, due to the low probability of high payouts, the cost of these 100 call options begins to fall for $T$ (and therefore liquidity) large enough; this is displayed in Figure 4(b).
(a) Fixed initial liquidity $L=\$100$.
(b) Varying initial liquidity $L=\$T$.
Figure 4: Section 6.2: Impact of placing a cap on the cost of 100 at-the-money call options with strike \$1.
Remark 11.
For both of our numerical experiments with puts and call options, we have only considered the case with a fixed strike price. We note that LBAMMs are designed to permit user to select any desired strike price. In such a way, even when considering digital puts and calls (to guarantee boundedness and eliminate the need for a cap, as was done for our European call option example above), we utilize a general probability space to allow users to freely set the strike price.